OpenStudy (anonymous):

@texaschic101 @ShadowLegendX

OpenStudy (texaschic101):

possibly..it kinda depends on what it is

OpenStudy (anonymous):

did you see it

OpenStudy (texaschic101):

your dealing with inequalities t + c <= 20 15t + 8c >= 150

OpenStudy (anonymous):

what??

OpenStudy (texaschic101):

the wording in the problem..." up to 20 hrs per week " and " at least 150 per week " indicates that these are inequalities

OpenStudy (anonymous):

t + c <= 20 15t + 8c >= 150

OpenStudy (anonymous):

^^ so those would be my inequalites?

OpenStudy (texaschic101):

I am not able to write the line that goes under the less then and greater then signs...so I just wrote an equal sign. yes...those are your inequalities

OpenStudy (anonymous):

ik, oh ok well thanks though I am not so confident with the inequalites....

OpenStudy (texaschic101):

they can be a bit confusing

OpenStudy (anonymous):

=D thanx

OpenStudy (texaschic101):

sure thing :)